OpenStudy (anonymous):

Given the function y= (kx-2)^3, where k is a constant, find the value of k such that y will have an inflection point at x= -1. Fully justify your answer.

OpenStudy (anonymous):

@jim_thompson5910 @zepdrix

jimthompson5910 (jim_thompson5910):

y= (kx-2)^3 y'= 3k(kx-2)^2 ... derive once y''= 6k^2(kx-2) ... derive again Now plug in x = -1 into the second derivative function and you should get y '' = 0, so you should now have 0 = 6k^2*(k*(-1) - 2) solve for k to get your answer(s)

OpenStudy (anonymous):

k= 0, -2?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

But since the problem says k is a nonzero constant, I should reject 0, so k is -2?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

i see "k is a constant", but if k is nonzero, then it's just k = -2

OpenStudy (anonymous):

Oh, oops I meant to write nonzero constant in the prob.

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

Thanks a bunch. Can I ask for further assistance?

jimthompson5910 (jim_thompson5910):

sure whats your question

OpenStudy (anonymous):

I have a few actually.

jimthompson5910 (jim_thompson5910):

ok how much is a few?

OpenStudy (anonymous):

umm, 3.

jimthompson5910 (jim_thompson5910):

ok go ahead

OpenStudy (anonymous):

A homeowner wishes to create an outdoor pet area surrounded by fencing, but comprised of two sections of equal area. The homeowner would like to enclose a total area of 150 square feet and wants to know the minimum amount of fencing to purchase and what dimensions to use. A layout of the pet area is shown below. Determine the total amount of fencing needed and the outer dimensions of the area.

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

since they have a common side (call it the width W) and equal area, they must share a common length L 

jimthompson5910 (jim_thompson5910):

Total Amount of Fencing: P = W+W+W + L+L+L+L P = 3W + 4L

jimthompson5910 (jim_thompson5910):

Each rectangle has an area of 75 square ft (cut 150 sq ft in half), so LW = 75 L = 75/W

jimthompson5910 (jim_thompson5910):

P = 3W + 4L P = 3W + 4(75/W) P = 3W + 300/W P = 3W^2/W + 300/W P = (3W^2 + 300)/W So the perimeter or the total length of fencing is (3W^2 + 300)/W ft

OpenStudy (anonymous):

Wait!

jimthompson5910 (jim_thompson5910):

what?

OpenStudy (anonymous):

I dont understand how you go from P=3W+300/W to P= 3W^2/W+ 300/W

jimthompson5910 (jim_thompson5910):

3W = 3W/1 3W = (3W/1)*(W/W) 3W = (3W^2)/W

OpenStudy (anonymous):

Oh, ok.